# HP-BERT: A framework for longitudinal study of Hinduphobia on social media via LLMs

### v1

@@ 5.1.1 Fine-tuning the HP-BERT Model
Table 7: Evaluation metrics of the HP-BERT model after fine-tuning on the SenWave dataset for sentiment analysis.
5.2 Data Analysis on Global COVID-19 X (Twitter) Dataset)
5.2.1 Analysis of COVID-19 Cases
@@
We first review the COVID-19 case trends from April 2020 to January 2022 across six countries, as shown in (Figure 4). We notice that India had a major peak in cases in April-May 2021, likely due to the Delta variant, while Brazil and Indonesia also saw significant peaks in mid-2021. The United Kingdom shows a rise toward late 2021, likely from the Omicron variant. Japan and Australia had relatively low case numbers throughout, with slight increases at the end of 2021. Overall, the trends reflect the varied impact of COVID-19 waves, and variants spread across these regions.
@@
Table 8: Tweet counts across six countries using HateBERT and HP-BERT on the global dataset and its subset.
5.2.2 Results of HateBERT model
We applied HateBERT [65] to analyze hate speech and abusive content during the COVID-19 pandemic across six countries, with the results summarised in (Table 8). To gain deeper insights into recurring themes within the discourse, we performed bigram and trigram analyses, as shown in (Figure 7 Panel-(a)) shows the most frequent bigrams, which reveal key patterns in the pandemic-related discourse. Bigrams such as ”free vaccine” and ”pvt hospital” highlight concerns around vaccine accessibility and healthcare disparities, while ”modi govt” and ”crore job” reflect criticisms of the Indian government and the economic challenges faced by citizens. (Figure 7 Panel-(b)) presents the most frequent trigrams, offering further insights into more complex interactions within the conversations. Trigrams such as ”daily wage labourer” and ”crore job loss” underscore the economic vulnerabilities faced by workers during the pandemic, while ”lockdown announced train” and ”stopped interstate travel” emphasise the disruptions caused by lockdown measures, particularly on mobility and commerce.
@@ 5.2.2 Results of HateBERT model
5.2.3 Results HP-BERT model
Using the HP-BERT model, we analyse Hindubhobic tweets and associated sentiments across countries and regions, providing insight into the perceptions of Hindu communities in various parts of the world during and after COVID-19. This can enable us to identify sentiment trends and potentially inform strategies for addressing harmful stereotypes and communal biases on social media platforms.
@@ 5.2.3 Results HP-BERT model
Finally, we applied our HP-BERT model to the Global COVID-19 X (Twitter) Dataset to analyze Hinduphobic tweets. (Figure 5) presents a multi-graph chart illustrating tweet counts across different countries, showing the monthly distribution of Hinduphobic tweet counts from April 2020 to January 2022. (Table 8) provide a detailed breakdown of the tweet counts across six countries where India recorded the highest number of Hinduphobic tweets, accounting for over 9,242 tweets, significantly surpassing other countries such as the United Kingdom, Australia, Japan, Indonesia, and Brazil, which reported much lower tweet counts. This higher number in India can be attributed to its large Hindu population [109], along with the presence of other significant religious communities such as Muslims and Christians. The United Kingdom ranked second, followed by Australia, Japan, Indonesia, and Brazil. This analysis highlights that Hinduphobic sentiment on Twitter during this period was predominantly concentrated in India. According to the 2011 Indian census [110], Hinduism encompassed 79.8% of the population followed by 14.2% Islam, and 2.3% Christianity.
We created an additional visualisation excluding India, as shown in Figure 6 to allow for a clearer view of Hinduphobic tweet counts across the remaining countries. We also present a distribution of Hinduphobic tweets (Figure 8) to review the fluctuating levels of Hinduphobic tweets, with notable peaks (April to May 2020 and April to May 2021). After these peaks, we can observe a general downward trend with periodic increases in March to April 2021, and September 2021. This distribution suggests varying intensities of Hinduphobic sentiment on Twitter throughout the observed period, with some months experiencing sharp rises in tweet volume while others show a noticeable decline.
@@
5.3 Analysis of HP-BERT results using HateBERT
Next, we apply HateBERT [65] to the results of HP-BERT to assess its effectiveness in detecting hate speech and abusive content, particularly in the context of Hinduphobia across six countries (Table 8). The detection results were comparatively lower because HateBERT is primarily trained on English language data, which may not effectively capture nuanced and culturally specific Hinduphobic rhetoric. HateBERT was pre-trained on a corpus of Reddit comments, making it more effective at detecting overtly hateful and abusive content. Furthermore, we generate bigrams and trigrams as shown in (Figure 11) to highlight recurring themes. These phrases reflect recurring themes that highlight narratives of Hinduphobia, particularly in the context of COVID-19 remedies, cow-related references, and religious or cultural stigmas. This analysis underscores the prevalence of stereotypes and how they manifest across different countries.
@@ 5.3 Analysis of HP-BERT results using HateBERT
5.4 HP-BERT Sentiment Analysis
We next employ the HP-BERT to perform sentiment analysis on Hinduphobic tweets, capturing a range of sentiment categories. (Table 7) provides evaluation metrics such as accuracy, precision, recall, and F1-score. Our analysis reveals a wide range of sentiments expressed in the tweets. We also examined user networks, activity patterns, and profile descriptions to identify common traits among users sharing Hinduphobic content. These insights offer a deeper understanding of the context and spread of Hinduphobia across social media platforms.
@@ 5.4 HP-BERT Sentiment Analysis
We next exclude the ”official report” category from the rest of the analysis to focus on the emotional content of the tweets for a country-wise examination to gain deeper insights into the sentiment distribution (Figure 16 Panel-(a)). The results show that ”annoyed” is the most dominant sentiment, with the highest percentage across all countries. The ”joking” sentiment consistently ranks second across all countries but appears at a significantly lower percentage compared to ”annoyed”. The ”optimistic” sentiment is also present at a lower level, however, in Japan, we observe nearly equal to ”annoyed”. The ”thankful” sentiment is scarcely present, with percentages approaching zero in all countries. (Figure 16 Panel-(b)) presents the distribution of ”empathetic”, ”pessimistic”, ”sad”, ”anxious” and ”denial” sentiments; where ”denial” emerges as the most frequently occurring sentiment across all countries. Other negative sentiments appear in minimal quantities across all countries. In summary, ”denial” reflects the prevailing negative sentiment against Hindus in the six countries, despite its relatively low frequency.
5.5 Polarity Score
We employ TextBlob and a custom weight strategy for polarity scores to quantify the sentiment expressed in text, categorising it as positive, negative, and neutral, where scores close to zero represent neutral sentiment. TextBlob [114] is a widely used Python library that provides various tools for NLP tasks, including sentiment analysis. The custom weight strategy involves assigning specific weights to each sentiment category (Table 5) to capture more complex sentiment expressions that traditional sentiment analysis tools may overlook. Figure-17 Panel-(a) presents the TextBlob polarity score revealing that a substantial portion of tweets has polarity scores close to zero, indicating a predominantly neutral sentiment. Furthermore, Figure 17 Panel-(b) presents the custom weight strategy, which demonstrates a concentration of scores near zero. However, the custom approach provides a more detailed representation, revealing multiple distinct peaks at various negative scores, notably at -0.20, which suggests more effectiveness in capturing subtle variations in sentiment.
@@ 5.5 Polarity Score
6 Discussion
@@
Our study analyses Hinduphobic tweets and their sentiment dynamics during the COVID-19 pandemic. We gained valuable insights into how public sentiment evolved during this period and how Hinduphobia manifested in different contexts for the selected countries. We first analysed the trends in COVID-19 cases (Figure 4) and their correlation with the prevalence of Hinduphobic tweets. As shown in (Figures 8) monthly distribution of Hinduphobic tweets and (Figure 5) country-wise distribution, peaks in COVID-19 cases often coincided with spikes in Hinduphobic tweets. This trend is further reflected in (Figure 18), which shows a drop in mean polarity scores, particularly during major outbreaks and significant public health interventions. These findings suggest a strong link between pandemic-related stressors and the rise in negative sentiment and Hinduphobic discourse.
@@
Our HP-BERT model predictions for Hinduphobic tweets, as presented in (Table 8), show the highest number of Hinduphobic tweets originating from India, followed by the United Kingdom, Australia, Japan and Brazil. For sentiment analysis (Figure 14), the emotions ”annoyed” and ”official report” were common, driven by tweets sharing anti-Hindu narratives, such as criticism of traditional practices, targeting Hindu festivals, or reacting to perceived bias in official reports. Our analysis highlighted a predominant presence of negative sentiments towards Hindus, with ”annoyed” and ”official report” being the most common emotions. These findings align with previous studies, such as Wang et al. [17] on Sinophobia sentiment analysis, and Chandra et al. [71, 105] on pandemic-related sentiment analysis, underscoring the robustness of sentiment analysis models in effectively capturing predominant emotions across diverse datasets. The analysis of the bigrams and trigrams in (Figure 9) for all countries and (Figure 10) for country-wise trends, revealed patterns in Hinduphobic tweets. These patterns are driven not only by the stereotypical association of Hindus with the coronavirus but also by a complex interplay of health, racial, and political factors. The frequent use of politically charged terms underscores how the discourse surrounding COVID-19 and its origins became highly politicised, with political narratives significantly influencing public sentiment during the pandemic.
@@
We also analysed the results of the HP-BERT model, as shown in (Table-8), and by using HateBERT [65] to assess its effectiveness in detecting hate speech and abusive content, particularly in the context of Hinduphobia across six countries. The detection results were lower for HateBERT because it is primarily trained on English-language Reddit data, which may not capture the nuanced, culturally specific Hinduphobic rhetoric. While HateBERT is more effective at identifying overtly hateful content, it struggles with the subtleties of Hinduphobia. Additionally, the bigrams and trigrams shown in (Figure 11) highlight recurring themes of Hinduphobia, particularly related to COVID-19 remedies, cow-related references, and religious stigmas, emphasising the persistence of stereotypes across countries. The analysis of sample tweets, as shown in Table 9 and (Table 10), indicates that the model performs well in accurately classifying sentiments in most cases. The samples generally align with expected sentiment patterns, showcasing the model’s effectiveness in capturing the predominant sentiments expressed in the tweets. However, specific examples reveal challenges in sentiment analysis, particularly those highlighted by Hussein [115], such as handling negation, sarcasm, and context-dependent language. For instance, consider the second tweet in (Table 9) ”Hindu Temples have donated huge sums of money so that people get food and all this is good”. This tweet was classified as ”Optimistic” with a positive polarity score of ”0.2,”. Similarly, in (Table 10), the tweet ”RT @NewsHandle: Religious events like Kumbh is reckless and spreading the virus. Why does India allow this?” was labelled as ”critical” and ”official report” with a polarity score of ”-0.3.” Although the classification of ”critical” aligns with the sentiment expressed, it fails to capture the underlying bias selectively targeting Hindu practices, highlighting the need for more context-sensitive sentiment analysis. These examples underscore the complexities of interpreting tweets with subtle or implicit biases. In the context of Hinduphobia, sarcastic, ironic, or biased language can easily evade detection or be misclassified. This highlights the importance of refining sentiment analysis models to address these challenges and ensure a more accurate understanding of the nuanced expressions of Hinduphobia in social media discourse.
@@
The dark historical practices associated with Hinduism, such as the caste system, sati (widow burning), child marriage, dowry, untouchability, temple entry restrictions, gender-based restrictions, animal sacrifice, the devadasi tradition, and caste hierarchy have long been reformed, abolished, and rarely practised in modern Hindu communities. The caste system [121], for instance, originated as a division of labour, but evolved into a rigid social hierarchy; however, it has been legally abolished by India’s Constitution and affirmative actions were introduced to promote equality. The practices such as sati [122] and child marriage [123] were outlawed in the 19th and early 20th centuries, and the dowry system [124] was criminalised in 1961. Temple entry [125] and gender-based restrictions [126] have largely been lifted due to reform movements, while animal sacrifice [127] and the devadasi system [128] have also seen significant cultural and legal rejection. Despite these advancements, Hindu society has often been misrepresented in Western media, which sometimes portrays these now-abolished practices as if they were active elements of Hinduism today [129, 130]. Western media still talks about outlawed practices as if they are part of Hinduism today [131, 132], without understanding the progress that has been made. This leads to a misleading view of Hinduism, showing it as stuck in the past, while ignoring the positive changes and efforts made to promote equality and social justice. By focusing on these old issues [133], the media overlooks the more inclusive and modern aspects of Hinduism, creating a negative image based on misunderstandings.

### v2

@@ V-A1 Fine-tuning the HP-BERT Model
TABLE VII: Evaluation metrics of the HP-BERT model after fine-tuning on the SenWave dataset for sentiment analysis.
V-B Comparative Model Analysis
We conducted a comprehensive evaluation against recent state-of-the-art Transformer-based models to validate the effectiveness of our HP-BERT model. We stress that HP-BERT incorporates multi-task capabilities (Hinduphobia detection and sentiment analysis), while baseline models focus on binary task classification. We designed our evaluation to ensure fair comparison while highlighting our model’s enhanced functionality. We conducted two distinct evaluation scenarios: (1) Baseline Model Comparison, where we directly downloaded pre-trained models from HuggingFace Hub and evaluated them on our ”Hinduphobic COVID-19 X (Twitter) Dataset” without any domain-specific adaptation, and (2) Fine-tuned Model Comparison, where we fine-tuned all baseline models using our ”Hinduphobic COVID-19 X (Twitter) Dataset” under identical conditions and hardware configurations used for the HP-BERT model.
We selected five baseline models representing diverse architectural innovations and training paradigms: RoBERTa [114] with optimized BERT training procedures and dynamic masking; DeBERTa [115] featuring disentangled attention mechanisms that separately encode content and position information; ELECTRA [116] employing replaced token detection objectives for efficient pre-training; DistilBERT [74] demonstrating knowledge distillation approaches with a parameter reduction of 60% while maintaining competitive performance; and HateBERT [76], a domain-specific model pre-trained on offensive content for the detection of hate speech.
The baseline models were evaluated using their default configurations from the HuggingFace Hub without any domain-specific adaptation. We accessed the models through the transformers.pipeline() API, employing sentiment analysis as a proxy for hate speech detection. All baseline models were subsequently fine-tuned on the ”Hinduphobic COVID-19 X (Twitter) Dataset” for the binary Hinduphobia detection task, using identical hyperparameters and training configurations as those employed for HP-BERT training. We accessed performance using the classification accuracy, precision, recall, and F1-score.
(Table VIII) demonstrates substantial performance disparities between HP-BERT and baseline model configurations. HP-BERT achieved 94.72% accuracy, significantly outperforming all baselines: HateBERT (67.21%), RoBERTa (66.45%), DeBERTa (65.18%), ELECTRA (63.92%), and DistilBERT (58.34%). The 27.51% performance gap over HateBERT, despite its hate speech specialization, highlights the critical importance of cultural and linguistic specificity in hate speech detection. These results validate our hypothesis that general-purpose models, even those trained for hate speech detection, do not perform well on culturally-specific tasks such as Hinduphobia detection.
Model Accuracy Precision Recall F1-Score
DistilBERT [74] 0.5834 0.5782 0.5886 0.5815
ELECTRA [116] 0.6392 0.6347 0.6436 0.6375
DeBERTa [115] 0.6518 0.6483 0.6552 0.6502
RoBERTa [114] 0.6645 0.6591 0.6698 0.6618
HateBERT [76] 0.6721 0.6507 0.6693 0.6802
HP-BERT (Ours) 0.9472 0.9474 0.9467 0.9465
TABLE VIII: Baseline model comparison on Hinduphobic tweets detection
When all models underwent identical fine-tuning on our dataset, the results shown in (Table IX) indicate that performance gaps reduced significantly but remained substantial, providing insights into both domain adaptation importance and architectural advantages. HP-BERT maintained superior performance (94.72% accuracy), followed by HateBERT (89.01%), RoBERTa (88.12%), DeBERTa (87.45%), ELECTRA (86.98%), and DistilBERT (84.23%). The convergence of baseline models to the 84-89% range demonstrates that domain-specific fine-tuning significantly improves performance across all architectures, while HP-BERT’s sustained advantage (5.71% over HateBERT, 6.60% over RoBERTa) indicates that our multi-stage fine-tuning approach and cultural context modeling provide architectural benefits beyond standard domain adaptation.
HP-BERT incorporates sentiment analysis capabilities through its multi-task learning approach, providing a significant advantage beyond classification performance. The multi-stage fine-tuning enables additional sentiment analysis capabilities without compromising primary task performance. This dual functionality allows comprehensive discourse analysis, as demonstrated in our longitudinal study of global Hinduphobic sentiment patterns, representing a significant advantage over single-task baseline models that require separate sentiment analysis pipelines.
Model Accuracy Precision Recall F1-Score
DistilBERT [74] 0.8423 0.8398 0.8447 0.8423
ELECTRA [116] 0.8698 0.8675 0.8721 0.8698
DeBERTa [115] 0.8745 0.8723 0.8767 0.8745
RoBERTa [114] 0.8812 0.8789 0.8835 0.8812
HateBERT [76] 0.8901 0.8878 0.8924 0.8901
HP-BERT (Ours) 0.9472 0.9474 0.9467 0.9465
TABLE IX: Fine-tuned model comparison on Hinduphobic tweets detection
V-C Data Analysis on Global COVID-19 X (Twitter) Dataset)
V-C1 Analysis of COVID-19 Cases
@@
We first review the COVID-19 case trends from April 2020 to January 2022 across six countries, as shown in (Figure 4). We notice that India had a major peak in cases in April- May 2021, likely due to the Delta variant, while Brazil and Indonesia also saw significant peaks in mid-2021. The United Kingdom shows a rise toward late 2021, likely from the Omicron variant. Japan and Australia had relatively low case numbers throughout, with slight increases at the end of 2021. Overall, the trends reflect the varied impact of COVID-19 waves and variants spread across these regions.
@@
TABLE X: Tweet counts across six countries using HateBERT and HP-BERT on the global dataset and its subset.
V-C2 Results of HateBERT model
We applied HateBERT [76] to analyze hate speech and abusive content during the COVID-19 pandemic across six countries, with the results summarised in (Table X). To gain deeper insights into recurring themes within the discourse, we performed bigram and trigram analyses, as shown in (Figure 7 Panel-(a)), which shows the most frequent bigrams, which reveal key patterns in the pandemic-related discourse. Bigrams such as ”free vaccine” and ”pvt hospital” highlight concerns around vaccine accessibility and healthcare disparities, while ”modi govt” and ”crore job” reflect criticisms of the Indian government and the economic challenges faced by citizens. (Figure 7 Panel-(b)) presents the most frequent trigrams, offering further insights into more complex interactions within the conversations. Trigrams such as ”daily wage labourer” and ”crore job loss” underscore the economic vulnerabilities faced by workers during the pandemic, while ”lockdown announced train” and ”stopped interstate travel” emphasise the disruptions caused by lockdown measures, particularly on mobility and commerce.
@@ V-C2 Results of HateBERT model
V-C3 Results HP-BERT model
Using the HP-BERT model, we analyze Hindubhobic tweets and associated sentiments across countries and regions, providing insight into the perceptions of Hindu communities in various parts of the world during and after the COVID-19 pandemic. This can enable us to identify sentiment trends and potentially inform strategies for addressing harmful stereotypes and communal biases on social media platforms.
@@ V-C3 Results HP-BERT model
Finally, we applied our HP-BERT model to the Global COVID-19 X (Twitter) Dataset to analyze Hinduphobic tweets. (Figure 5) presents a multi-graph chart illustrating tweet counts across different countries, showing the monthly distribution of Hinduphobic tweet counts from April 2020 to January 2022. Table X provides a detailed breakdown of the tweet counts across six countries, where India recorded the highest number of Hinduphobic tweets, accounting for over 9,242 tweets, significantly surpassing other countries such as the United Kingdom, Australia, Japan, Indonesia, and Brazil, which reported much lower tweet counts. This higher number in India can be attributed to its large Hindu population [117], along with the presence of other significant religious communities, such as Muslims and Christians. The United Kingdom ranked second, followed by Australia, Japan, Indonesia, and Brazil. This analysis highlights that the Hinduphobic sentiment on Twitter during this period was predominantly concentrated in India. According to the 2011 Indian census [118], Hinduism encompassed 79.8% of the population, followed by 14.2% Islam and 2.3% Christianity.
We created an additional visualization excluding India, as shown in (Figure 6) to allow a clearer view of Hinduphobic tweet counts in the remaining countries. We also present a distribution of Hinduphobic tweets (Figure 8) to review the fluctuating levels of Hinduphobic tweets, with notable peaks (April to May 2020 and April to May 2021). After these peaks, we can observe a general downward trend with periodic increases from March to April 2021 and September 2021. This distribution suggests varying intensities of Hinduphobic sentiment on Twitter throughout the observed period, with some months experiencing sharp rises in tweet volume while others show a noticeable decline.
@@
V-D Analysis of HP-BERT results using HateBERT
Next, we apply HateBERT [76] to the results of HP-BERT to assess its effectiveness in detecting hate speech and abusive content, particularly in the context of Hinduphobia across six countries (Table X). The detection results were comparatively lower because HateBERT is primarily trained on English language data, which may not effectively capture nuanced and culturally specific Hinduphobic rhetoric. HateBERT was pre-trained on a corpus of Reddit comments, making it more effective at detecting overtly hateful and abusive content. Furthermore, we generate bigrams and trigrams as shown in (Figure 11) to highlight recurring themes. These phrases reflect recurring themes that highlight narratives of Hinduphobia, particularly in the context of COVID-19 remedies, cow-related references, and religious or cultural stigmas. This analysis underscores the prevalence of stereotypes and how they manifest across different countries.
@@ V-D Analysis of HP-BERT results using HateBERT
V-E Statistical Significance of Temporal and Geographical Patterns
Our statistical analysis of the results confirms significant variations in Hinduphobic discourse across both temporal and geographical dimensions. The cross-country comparison using one-way Analysis of Variance (ANOVA) revealed significant differences in tweet volumes (F(5,156) = 18.67, p ¡ 0.001), with India accounting for 84.7% of total Hinduphobic content (9,242 tweets), significantly higher than the UK (746 tweets), Australia (479 tweets), Japan (389 tweets), Brazil (352 tweets), and Indonesia (356 tweets) as shown in (Table X) (Tukey’s Honestly Significant Difference (HSD), all p ¡ 0.001).
Temporal analysis using the Kruskal-Wallis test showed significant differences across pandemic phases (H = 12.45, p = 0.002). The during-COVID period (April 2020-December 2021) exhibited significantly higher tweet volumes (Mean = 142.3 tweets/month, Standard Deviation (SD) = 89.7) compared to pre-COVID baseline (January-March 2020: M = 67.2 tweets/month, SD = 34.5) and the post-peak period (January 2022 onwards: Mean = 98.4 tweets/month, SD = 52.1), as illustrated in (Figure 4) and (Figure 5). The results of our linear regression model revealed a significant increase trend during the study period for India ($\beta$ = 3.42, p = 0.018, R² = 0.23) and the UK ($\beta$ = 1.87, p = 0.041, R² = 0.16), confirmed by Mann-Kendall trend analysis (India: $\tau$ = 0.234, p = 0.012; UK: $\tau$ = 0.189, p = 0.038).
Correlation analysis between monthly COVID-19 case counts (Figure 8) and Hinduphobic content revealed moderate positive associations across countries: India (r = 0.428, p = 0.016), UK (r = 0.391, p = 0.029), Australia (r = 0.312, p = 0.087), with weaker correlations in other countries. The peak coincidence analysis showed that 68% of the major spikes in Hinduphobic content (defined as¿ 2 SD above mean) occurred within 4-6 weeks of the COVID-19 surge periods, significantly higher than expected by chance ($\chi^{2}$ = 6.73, p = 0.034). These findings, supported by the bigram and trigram analyses (Figure 9) and (Figure 10), provide robust statistical evidence for pandemic-driven amplification of anti-Hindu sentiment, particularly in countries with larger Hindu diaspora populations.
V-F HP-BERT Sentiment Analysis
We next employ the HP-BERT to perform sentiment analysis on Hinduphobic tweets, capturing a range of sentiment categories. (Table VII) provides evaluation metrics such as accuracy, precision, recall, and F1-score. Our analysis reveals a wide range of sentiments expressed in the tweets. We also examined user networks, activity patterns, and profile descriptions to identify common traits among users sharing Hinduphobic content. These insights offer a deeper understanding of the context and spread of Hinduphobia across social media platforms.
@@ V-F HP-BERT Sentiment Analysis
We next exclude the ”official report” category from the rest of the analysis to focus on the emotional content of the tweets for a country-wise examination to gain deeper insights into the sentiment distribution (Figure 16 Panel-(a)). The results show that ”annoyed” is the most dominant sentiment, with the highest percentage across all countries. The ”joking” sentiment consistently ranks second across all countries, but appears at a significantly lower percentage compared to ”annoyed”. The ”optimistic” sentiment is also present at a lower level; however, in Japan, we observe nearly equal amounts of ”annoyed”. The ”thankful” sentiment is scarcely present, with percentages approaching zero in all countries. (Figure 16 Panel-(b)) presents the distribution of ”empathetic”, ”pessimistic”, ”sad”, ”anxious” and ”denial” sentiments, where ”denial” emerges as the most frequently occurring sentiment across all countries. Other negative sentiments appear in minimal quantities across all countries. In summary, ”denial” reflects the prevailing negative sentiment against Hindus in the six countries despite its relatively low frequency.
V-G Polarity Score
We employ TextBlob and a custom weight strategy for polarity scores to quantify the sentiment expressed in text, categorizing it as positive, negative, and neutral, where scores close to zero represent neutral sentiment. TextBlob [122] is a widely used Python library that provides various tools for NLP tasks, including sentiment analysis. The custom weight strategy involves assigning specific weights to each sentiment category (Table V) to capture more complex sentiment expressions that traditional sentiment analysis tools may overlook. (Figure 17 Panel-(a)) presents the TextBlob polarity score, revealing that a substantial portion of tweets has polarity scores close to zero, indicating a predominantly neutral sentiment. Furthermore, (Figure 17 Panel-(b)) presents the custom weight strategy, which demonstrates a concentration of scores near zero. However, the custom approach provides a more detailed representation, revealing multiple distinct peaks at various negative scores, notably at -0.20, which suggests more effectiveness in capturing subtle variations in sentiment.
@@ V-G Polarity Score
VI Discussion
Our study presents the first comprehensive computational framework for analyzing Hinduphobia during a COVID-19 crisis, making four primary contributions to understanding religious discrimination on social media. First, the creation of the first ”Hinduphobic COVID-19 X (Twitter) Dataset” with 8,000 rigorously annotated tweets through human-in-the-loop annotation and manual verification. Second, we developed HP-BERT, achieving 94.72% accuracy in detecting Hinduphobic content, and its sentimental analysis significantly outperformed five state-of-the-art transformer models, with improvements ranging from 27.51% to 36.38% over baseline applications. Third, we established statistically significant correlations (r = 0.312-0.428) between COVID-19 case surges and Hinduphobic discourse volume across six countries, providing the first quantitative evidence of pandemic-driven religious discrimination amplification. Fourth, we conducted a longitudinal analysis of 27.4 million tweets, revealing previously undocumented geographical and temporal patterns of anti-Hindu sentiment during COVID-19.
The correlation analysis between COVID-19 cases (Figure 4) and Hinduphobic tweet prevalence reveals compelling evidence of crisis-driven discrimination. Our statistical analysis using one-way ANOVA showed significant differences in tweet volumes across countries (F(5,156) = 18.67, p ¡ 0.001), with India accounting for 84.7% of detected content (9,242 tweets). Peak coincidence analysis demonstrated that 68% of major Hinduphobic content spikes (defined as ¿2 SD above mean) occurred within 4-6 weeks of COVID-19 surge periods, significantly higher than expected by chance ($X^{2}$ = 6.73, p = 0.034). The temporal lag suggests that pandemic stress creates conditions where minority communities become targets for blame and discrimination.
Our evaluation against five Transformer models validates HP-BERT’s effectiveness for culturally-specific hate speech detection. The baseline comparison revealed outperforming existing models by significant margins with improvements of +27.51% to +36.38% over baseline models (Table VIII) and +5.71% to +10.49% over fine-tuned (Table IX). The sentiment analysis results reveal concerning patterns absent in general pandemic discourse. Unlike studies showing empathetic responses in COVID-19 discussions, Hinduphobic tweets exhibited a stark absence of empathy, with ”annoyed” (9,000+ tweets) and ”official report” dominating sentiment categories. The bigram and trigram analysis (Figure 9) and (Figure 10) reveals how traditional stereotypes resurface through contemporary health narratives. The phrases such as ”cow urine cure” and ”Hindu superspreaders” represent continuations of colonial-era caricatures that have been amplified through social media echo chambers during crisis periods.
@@ VI Discussion
Our study analyses Hinduphobic tweets and their sentiment dynamics during the COVID-19 pandemic. We gained valuable insights into how public sentiment evolved during this period and how Hinduphobia manifested in different contexts for the selected countries. We first analysed the trends in COVID-19 cases (Figure 4) and their correlation with the prevalence of Hinduphobic tweets. As shown in (Figure 8) monthly distribution of Hinduphobic tweets and (Figure 5) country-wise distribution, peaks in COVID-19 cases often coincided with spikes in Hinduphobic tweets. This trend is further reflected in (Figure 18), which shows a drop in mean polarity scores, particularly during major outbreaks and significant public health interventions. These findings suggest a strong link between pandemic-related stressors and the rise in negative sentiment and Hinduphobic discourse.
@@
Our HP-BERT model predictions for Hinduphobic tweets, as presented in (Table X), show the highest number of Hinduphobic tweets originating from India, followed by the United Kingdom, Australia, Japan and Brazil. For sentiment analysis (Figure 14), the emotions ”annoyed” and ”official report” were common, driven by tweets sharing anti-Hindu narratives, such as criticism of traditional practices, targeting Hindu festivals, or reacting to perceived bias in official reports. Our analysis highlighted a predominant presence of negative sentiments towards Hindus, with ”annoyed” and ”official report” being the most common emotions. These findings align with previous studies, such as Wang et al. [17] on Sinophobia sentiment analysis, and Chandra et al. [82, 109] on pandemic-related sentiment analysis, underscoring the robustness of sentiment analysis models in effectively capturing predominant emotions across diverse datasets. The analysis of the bigrams and trigrams in (Figure 9) for all countries and (Figure 10) for country-wise trends revealed patterns in Hinduphobic tweets. These patterns are driven not only by the stereotypical association of Hindus with the coronavirus but also by a complex interplay of health, racial, and political factors. The frequent use of politically charged terms underscores how the discourse surrounding COVID-19 and its origins became highly politicised, with political narratives significantly influencing public sentiment during the pandemic.
@@
We also analysed the results of the HP-BERT model, as shown in (Table X), and by using HateBERT [76] to assess its effectiveness in detecting hate speech and abusive content, particularly in the context of Hinduphobia across six countries. The detection results were lower for HateBERT because it is primarily trained on English-language Reddit data, which may not capture the nuanced, culturally specific Hinduphobic rhetoric. While HateBERT is more effective at identifying overtly hateful content, it struggles with the subtleties of Hinduphobia. Additionally, the bigrams and trigrams shown in (Figure 11) highlight recurring themes of Hinduphobia, particularly related to COVID-19 remedies, cow-related references, and religious stigmas, emphasising the persistence of stereotypes across countries. The analysis of sample tweets, as shown in (Table XI) and (Table XII), indicates that the model performs well in accurately classifying sentiments in most cases. The samples generally align with expected sentiment patterns, showcasing the model’s effectiveness in capturing the predominant sentiments expressed in the tweets. However, specific examples reveal challenges in sentiment analysis, particularly those highlighted by Hussein [123], such as handling negation, sarcasm, and context-dependent language. For instance, consider the second tweet in (Table XI) ”Hindu Temples have donated huge sums of money so that people get food and all this is good”. This tweet was classified as ”Optimistic” with a positive polarity score of ”0.2,”. Similarly, in (Table XII), the tweet ”RT @NewsHandle: Religious events like Kumbh is reckless and spreading the virus. Why does India allow this?” was labelled as ”critical” and ”official report” with a polarity score of ”-0.3.” Although the classification of ”critical” aligns with the sentiment expressed, it fails to capture the underlying bias selectively targeting Hindu practices, highlighting the need for more context-sensitive sentiment analysis. These examples underscore the complexities of interpreting tweets with subtle or implicit biases. In the context of Hinduphobia, sarcastic, ironic, or biased language can easily evade detection or be misclassified. This highlights the importance of refining sentiment analysis models to address these challenges and ensure a more accurate understanding of the nuanced expressions of Hinduphobia in social media discourse.
@@
The dark historical practices associated with Hinduism, such as the caste system, sati (widow burning), child marriage, dowry, untouchability, temple entry restrictions, gender-based restrictions, animal sacrifice, the devadasi tradition, and caste hierarchy have long been reformed, abolished, and rarely practised in modern Hindu communities. The caste system [129], for instance, originated as a division of labour but evolved into a rigid social hierarchy; however, it has been legally abolished by India’s Constitution and affirmative actions were introduced to promote equality. Practices such as sati [130] and child marriage [131] were outlawed in the 19th and early 20th centuries, and the dowry system [132] was criminalised in 1961. Temple entry [133] and gender-based restrictions [134] have largely been lifted due to reform movements, while animal sacrifice [135] and the devadasi system [136] have also seen significant cultural and legal rejection. Despite these advancements, Hindu society has often been misrepresented in Western media, which sometimes portrays these now-abolished practices as if they were active elements of Hinduism today [137, 138]. Western media still talks about outlawed practices as if they are part of Hinduism today [139, 140], without understanding the progress that has been made. This leads to a misleading view of Hinduism, showing it as stuck in the past while ignoring the positive changes and efforts made to promote equality and social justice. By focusing on these old issues [141], the media overlooks the more inclusive and modern aspects of Hinduism, creating a negative image based on misunderstandings.